# Model Equality Testing: Which Model Is This API Serving?

### v1

@@ Abstract
Users often interact with large language models through black-box inference APIs, both for closed- and open-weight models (e.g., Llama models are popularly accessed via Amazon Bedrock and Azure AI Studios). In order to cut costs or add functionality, API providers may quantize, watermark, or finetune the underlying model, changing the output distribution — often without notifying users. We formalize detecting such distortions as Model Equality Testing, a two-sample testing problem, where the user collects samples from the API and a reference distribution, and conducts a statistical test to see if the two distributions are the same. We find that tests based on the Maximum Mean Discrepancy between distributions are powerful for this task: a test built on a simple string kernel achieves a median of 77.4% power against a range of distortions, using an average of just 10 samples per prompt. We then apply this test to commercial inference APIs for four Llama models, finding that 11 out of 31 endpoints serve different distributions than reference weights released by Meta.
@@ 1 Introduction
Since running a large language model requires compute and technical expertise, many users rely on black-box APIs to handle inference. This reliance applies to both closed-weight models, like GPT and Claude, and open-weight ones: for example, Amazon Bedrock, Microsoft Azure, and the seven other companies in Figure 1 all compete to offer Llama models as a service. While users can sample from black-box APIs, they have little to no insight into the underlying implementation of the model, including into questions like:
@@
    How has the API modified the language model’s distribution? To drive down costs, API providers may quantize or prune large model weights; they may also watermark outputs or incorrectly implement some decoding parameters. These changes distort the resulting distribution of completions. The problem is when such distortions are undisclosed: users assume that calling a third-party API is exactly equivalent to working with the original model. For example, benchmarks like HELM (Liang et al., 2022) evaluate models through third-party APIs, but quantized or watermarked models may be less capable than the intended model. Can we test if an API’s distribution differs from reference weights, and if so, by how much?
@@
    Is the API changing over time? Language model inference endpoints may also drift over time without notifying users (Chen et al., 2023; Eyuboglu et al., 2024), e.g., due to finetuning or updates to the inference stack. Unstable APIs affect research reproducibility (Pozzobon et al., 2023) and can disrupt user productivity in human-AI teams (Bansal et al., 2019). Can we track if a black-box API’s distribution has changed over time?
@@
These concerns are important to address: tens of thousands of developers already rely on black-box inference APIs for applications (Amazon, 2024), and this dependence will increase as LLMs — and the corresponding infrastructure costs for hosting — grow larger. For example, most users must rely on third-party APIs to use Llama 3.1 405B because of its size.
@@
The current approach to this problem is for an outside auditor to monitor APIs’ accuracies on multiple-choice or short-answer benchmarks (ArtificialAnalysis, 2024); these studies typically decode from the language model greedily. Such audits can be a poor match for a particular user’s needs. Greedy decoding only checks that the modes of the next-token distributions match, rather than the overall distribution over completions. This limitation is problematic because users often sample from models. Further, short-answer benchmarks cover only a small slice of possible prefixes, which may significantly differ from users’ applications: common applications like code generation, dialogue, and summarization are longform tasks. Ideally, users would be able to personally audit APIs on their custom applications. Such a method should be sample-efficient, apply to tasks without clear evaluation metrics, and assess with confidence whether the overall distribution of completions has shifted in a statistically significant way.
@@
Figure 1: (Left) We formalize auditing black-box language model APIs as Model Equality Testing, a two-sample hypothesis testing problem. This enables us to answer questions such as an API’s faithfulness to a reference distribution and stability over time. (Right) We evaluate candidate tests and select one based on the Maximum Mean Discrepancy between distributions (Gretton et al., 2012). We then apply this method to commercial inference APIs serving Llama models, finding that 11 out of 31 endpoints deviate from reference weights released by Meta.
@@
We provide such a method. Suppose a user wants to audit a an API on their task of interest. The user collects two samples: one from a reference distribution $P$ and one from the test API’s distribution $Q$. For example, to answer if an API has modified the distribution of an open-weight model, $P$ might be from reference model weights released on Hugging Face. To answer if the API is changing over time, $P$ might be from the API at an earlier point in time (Figure 1 left). The user then conducts a two-sample test for whether $P=Q$ to examine if the API’s distribution is statistically indistinguishable from the reference.
@@
Our setting is challenging because the distributions being compared are high-dimensional: they are defined over multi-token completions from large vocabularies. We show that two-sample kernel tests based on estimating the Maximum Mean Discrepancy (MMD) between $P$ and $Q$ (Gretton et al., 2012) are powerful tools for this setting, as kernel tests allow us to specify a featurization that reduces dimensionality. We find that a simple string kernel based on the Hamming distance between completions works particularly well with few samples. In simulations (§4), this test achieves a median of 77.4% power against a wide range of distortions — e.g., quantization, watermarking, and finetuning — using an average of just 10 samples per prompt for distributions over 20–25 prompts. We then apply this test to nine commercial inference API providers across four Llama models (§5, Figure 1 right). Our test flags 11 out of these 31 endpoints.
@@
Because our test statistic is an estimate of a distance, we also explore how this same machinery can be used to quantify statistical distances between black-box endpoints. In §4.3, we estimate pairwise distances between the output distributions of 13 different language models — without requiring log probability access — and find that models within the same family (e.g., the Llama family, GPT-3.5 family, or GPT-4o family) output more similar distributions than models within the same size range (e.g., 7B or 70B models). In §5, we apply this method to estimate the effect size of deviations between API endpoints and reference weights, finding that some deviations are quite serious: some APIs’ implementations are further from the reference weights than if the provider had substituted in an entirely different language model.
@@
In summary, our contributions are as follows: we unify several API auditing tasks under the formalization of Model Equality Testing, a two-sample distribution testing problem, and extensively empirically validate kernel-based tests for this problem. We then apply this test to audit popular commercial inference APIs. To enable users to audit APIs for custom applications, we open-source a Python package. We also aim to encourage future research in Model Equality Testing by releasing a dataset of 1 million LLM completions from five models. [^1] Package, experiment code, and dataset: https://github.com/i-gao/model-equality-testing.
@@ 2 The Model Equality Testing problem
Suppose an auditor is interested in a task parameterized by a distribution $\pi$ over $m$ prompts and a maximum completion length of $L$ tokens. The auditor has sample access to a reference distribution $P$ and API distribution $Q$, both operating on the same vocabulary $\mathcal{V}$ and with the same decoding parameters. The auditor samples $N$ prompt-completion pairs [^2] In Appendix C.4, we discuss how to extend this setup to unequal sample sizes. $z:=(x,y)$ from each distribution:
@@ 3 Method
$\widehat{\text{MMD}}(\mathcal{D}_{Q},\mathcal{D}_{P})$ $=\frac{1}{N(N-1)}\left[\sum_{z,z^{\prime}\in\mathcal{D}_{P}}k(z,z^{\prime})+\sum_{z,z^{\prime}\in\mathcal{D}_{Q}}k(z,z^{\prime})\right]-\frac{2}{N^{2}}\sum_{z\in\mathcal{D}_{P}}\sum_{z^{\prime}\in\mathcal{D}_{Q}}k(z,z^{\prime}).$ (4)
@@
The choice of kernel $(k,\phi)$ determines the test’s semantics and power. For example, setting $\phi(z)$ to be an indicator of whether $y$ is an accurate completion for $x$ leads to rejecting the null when $P,Q$ result in substantially different task accuracies. However, because $\phi$ is relatively coarse, the MMD may be zero even when $P$ and $Q$ are quite different, limiting the test’s power. At the other extreme are universal kernels, which guarantee that the MMD is zero if and only if $P=Q$ (Gretton et al., 2012). One such universal kernel for strings is the computationally expensive all-substrings kernel (Borgwardt et al., 2006):
@@
where $y$ shorter than $L$ is right-padded with a special token. Intuitively, a test based on this kernel rejects if a significantly larger number of substitutions are needed to align completions between $\mathcal{D}_{P}$ and $\mathcal{D}_{Q}$ than within each sample. The associated Hamming MMD is a pseudo-metric, as it is zero when $P=Q$ and obeys the triangle inequality, but may not separate all distributions (Appendix A.1). Despite this limitation, we find in the following sections that this kernel is empirically effective and well-suited to common distortions we encounter with language models: quantization, watermarking, finetuning, and related distortions tend to result in large inter-sample Hamming distances.
@@ 4.1 Detecting quantization and watermarking
In this section, the reference distribution $P$ represents weights published on Hugging Face inferenced at full precision. We evaluate if tests can distinguish $P$ from alternative distributions $Q$:
@@
To test generalization across models, we repeat evaluations for 5 instruction-tuned models: Mistral 7B Instruct (Jiang et al., 2023), Llama-3 8B and 70B Instruct, and Llama-3.1 8B and 70B Instruct (Meta, 2024).
Tests. We compare three choices of kernels: the Hamming kernel (7), the all-substrings kernel (5), and the one-hot kernel (6). Additionally, we compare two tests from the multinomial testing literature (Balakrishnan & Wasserman, 2018; Bhattacharya & Valiant, 2015):
@@
Figure 2: (Left) Sample complexity of tests. At an average of just 10 samples per prompt, the Hamming MMD test is able to detect quantization and watermarking with nontrivial power. Curves first median power across alternative distributions $Q$, averaged over language models and prompt distributions, with shaded standard errors. Results stratified by language model and alternative are in Appendix C. (Middle) While other tests rapidly degrade in power when the user is interested in longer completions, the Hamming MMD test maintains power best across completion lengths. (Right) Power of the Hamming MMD test, stratified by alternative distribution. The test is significantly less powerful against the fp16 alternative.
@@ 4.3 Distinguishing if samples come from different language models
We next explore if the Hamming MMD test can distinguish whether two bodies of text are generated from the same language model. In this setting, $P$ and $Q$ are two different language models, drawn from a pool of 13 instruction-tuned models (Figure 3 right), including eight open-weight models (Abdin et al., 2024; Groeneveld et al., 2024; Team et al., 2024) and and five OpenAI closed-weight models. In order to compare models with different tokenizers, samples must be compared in character space. We sample $L=50$-token completions to the Wikipedia language modeling task as before, and then we decode these to characters, ignoring special tokens. The vocabulary of interest is now all Unicode characters ($|\mathcal{V}|\approx$ 1.1M), the maximum completion length is $L\approx$1000$$ characters, and we test with $N=10m$ samples.
@@
Llama-3 8B 0.76
Llama-3 70B 1.00
Llama-3.1 70B 0.75
OLMo 7B 1.00
gpt-4o-mini 1.00
Llama-3.1 8B (watermark) 0.32
Llama-3.1 8B (int8) 0.07
Llama-3.1 8B (nf4) 1.00
Figure 3: (Upper left) The Hamming MMD test is able to detect when Llama-3 8B has been finetuned on datasets of $1000$ samples, even after a single epoch. Power is higher, earlier, when the finetuning distribution is i.i.d. with the testing distribution. (Lower left) The Hamming MMD test can also detect when two models are different with near-perfect power. Full results are in Appendix C. (Right) The MMD framework allows us to estimate statistical distance between any models from which we can draw samples. The cells show average estimated MMDs over $10$ bootstraps. All standard errors were less than 0.001. Rows are sorted using spectral clustering with two components. Models within a family are typically clustered together, suggesting that factors like training data, rather than scale, determine model similarity.
We find that all model swaps are detectable with 100% power, except for pairs within the Llama family, e.g., Llama-3.1 8B and Llama-3.1 70B (75% power) or Llama-3 8B and Llama-3.1 8B (76% power; Figure 3 lower left). Model swaps are significantly more detectable than watermarking or quantization. In Appendix D.2, we show qualitative examples of completions from different models.
@@
A useful feature of the MMD tests is that the test statistic is an estimator of a distance. As a result, we can reuse the machinery to quantify the degree to which two models differ by estimating $\mathbb{E}_{x\sim\pi}[\text{MMD}(P(\cdot\mid x),Q(\cdot\mid x))]$. [^4] Note that the Hamming MMD is a pseudometric: $\text{MMD}_{\text{Hamming}}(P,Q)=0$ does not imply $P=Q$ (Appendix A.1). Figure 3 (right) estimates the Hamming MMD between all pairs of models. To decrease estimator error, we increase the sample size to $N=100m$ samples from each model and report the average $\widehat{\text{MMD}}$ over $10$ simulations, along with standard errors. We observe that models within a family are typically clustered together, suggesting that training data, rather than scale, determines model similarity. Surprisingly, while generations of Llama models (3 and 3.1) are close in distance, some generations of GPT models (e.g., 4-preview and 3.5-turbo) are not. This result suggests the use of significantly different training data or procedures between these models.
@@ 5 Auditing inference API providers
As a case study, we now apply our test to 31 commercial inference endpoints for Meta’s Llama models (Figure 1). We seek to understand whether these APIs have modified the distribution of the underlying model. The 31 inference endpoints are distributed across four Llama models (instruction-tuned versions of Llama-3 8B and 70B, Llama-3.1 8B and Llama-3.1 70B) and nine providers: Amazon Bedrock, Anyscale, [^5] We collected samples from Anyscale’s serverless endpoints from before they were deprecated in August 2024. Azure AI Studio, DeepInfra, Fireworks AI, Groq, Perplexity, Replicate, and Together.ai. Note that at the time of writing, only two companies publicize using distribution-altering optimizations: Fireworks AI notes that they employ semantic caching, and Together notes the use of quantization. [^6] Sources: Fireworks AI home page, Together.ai product page
We are interested in whether APIs are indistinguishable from the model at commonly accepted precisions. Specifically, we consider two possible null distributions, both derived from model weights released by Meta on Hugging Face: $P_{1}$ is the full-precision model weights, and $P_{2}$ is the fp16-precision model weights. Our aim is to distinguish between the hypotheses

### v2

@@ Abstract
Users often interact with large language models through black-box inference APIs, both for closed- and open-weight models (e.g., Llama models are popularly accessed via Amazon Bedrock and Azure AI Studios). In order to cut costs or add functionality, API providers may quantize, watermark, or finetune the underlying model, changing the output distribution — possibly without notifying users. We formalize detecting such distortions as Model Equality Testing, a two-sample testing problem, where the user collects samples from the API and a reference distribution, and conducts a statistical test to see if the two distributions are the same. We find that tests based on the Maximum Mean Discrepancy between distributions are powerful for this task: a test built on a simple string kernel achieves a median of 77.4% power against a range of distortions, using an average of just 10 samples per prompt. We then apply this test to commercial inference APIs from Summer 2024 for four Llama models, finding that 11 out of 31 endpoints serve different distributions than reference weights released by Meta.
@@ 1 Introduction
Since running a large language model requires compute and technical expertise, many users rely on black-box APIs to handle inference. This applies to both closed-weight models, like GPT and Claude, and open-weight ones: e.g., Amazon Bedrock, Microsoft Azure, and the seven other companies in Figure 1 all compete to offer Llama models as a service. While users can sample from black-box APIs, they have little to no insight into the underlying implementation of the model, including into questions like:
@@
    How has the API modified the language model’s distribution? To drive down costs, API providers may quantize or prune large model weights; they may also watermark outputs or incorrectly implement some decoding parameters. These changes distort the resulting distribution of completions. The problem is when such distortions are undisclosed: users assume that calling a third-party API is exactly equivalent to working with the original model. For example, benchmarks like HELM (Liang et al., 2022) evaluate models through third-party APIs, but quantized or watermarked models may be less capable than the intended model.
@@
    Is the API changing over time? Language model inference endpoints may also drift over time without notifying users (Chen et al., 2023; Eyuboglu et al., 2024), e.g., due to finetuning or updates to the inference stack. Unstable APIs affect research reproducibility (Pozzobon et al., 2023) and can disrupt user productivity in human-AI teams (Bansal et al., 2019).
@@
Under the status quo, neither users nor regulators have a way to rigorously answer these questions for themselves.111 These problems are already experienced by users: e.g., see threads 1, 2 and 3. These concerns are important to address: tens of thousands of developers already rely on black-box inference APIs for applications (Amazon, 2024), and this dependence will increase as LLMs — and the corresponding infrastructure costs for hosting — grow larger. For example, most users must rely on third-party APIs to use Llama 3.1 405B because of its size.
@@
The current approach to this problem is for an outside auditor to monitor APIs’ accuracies on multiple-choice or short-answer benchmarks (ArtificialAnalysis, 2024); these studies typically decode from the language model greedily. Such audits can be a poor match for user needs. Greedy decoding only checks that the modes of the next-token distributions match, rather than the overall distribution over completions, which is problematic because users often sample from models. Short-answer benchmarks cover only a small slice of possible prefixes, which may significantly differ from a particular user’s task: e.g., common applications like code generation, dialogue, and summarization are longform tasks. Ideally, users could personally audit APIs on their custom tasks. Such a method should be sample-efficient, apply to tasks without automated verifiers, and assess with confidence whether the overall distribution of completions has shifted in a statistically significant way.
@@
Figure 1: (Left) We formalize auditing black-box language model inference APIs as Model Equality Testing. This enables us to assess an API’s faithfulness to a reference distribution and its stability over time. (Right) We evaluate candidate tests and apply the most powerful one to Llama model APIs from Summer 2024, finding that 11 of 31 endpoints deviate from reference weights released by Meta.
@@
We provide such a method. Suppose a user wishes to audit a an API on their task of interest. The user collects two samples: one from a reference distribution $P$ and one from the test API’s distribution $Q$. For example, to answer if an API has modified the distribution of an open-weight model, $P$ might be from reference model weights released on Hugging Face. To answer if the API is changing over time, $P$ might be from the API at an earlier point in time (Figure 1 left). The user then conducts a two-sample test for whether $P=Q$ to examine if the API’s distribution is statistically indistinguishable from the reference.
@@
Our setting is challenging because the distributions being compared are high-dimensional: they are defined over multi-token completions from large vocabularies. Two-sample kernel tests based on estimating the Maximum Mean Discrepancy (MMD) between $P$ and $Q$ (Gretton et al., 2012) are flexible tools for this setting, as they allow us to specify a featurization to reduce dimensionality. We find that a simple string kernel based on the Hamming distance between completions is particularly sample-efficient. In simulations (§4), this test achieves a median of 77.4% power against a wide range of distortions — e.g., quantization, watermarking, and finetuning — using an average of just 10 samples per prompt for distributions over 20–25 prompts. We then apply this test to nine commercial inference API providers across four Llama models (§5, Figure 1 right). Our test flags 11 out of these 31 endpoints, with each audit costing less than \$1.
@@
Because our test statistic is an estimate of a distance, we also explore how this machinery can quantify statistical distances between black-box endpoints. In §4.3, we estimate pairwise distances between the output distributions of 13 language models — without requiring log probability access — and find that models within the same family (e.g., the Llama family or GPT-3.5 family) output more similar distributions than models within the same size range (e.g., 7B or 70B models). In §5, we estimate the effect size of deviations between API endpoints and reference weights, finding that some implementations are further from the reference weights than if the provider had substituted in an entirely different language model.
Summary of contributions.
@@
We unify several API auditing tasks under the formalization of Model Equality Testing, a two-sample distribution testing problem, and empirically validate kernel-based tests for this problem. We then apply this test to audit popular commercial inference APIs. To enable users to audit APIs for custom applications, we open-source a Python package. We also encourage future research in Model Equality Testing by releasing a dataset of 1 million LLM completions from five models. [^2] Package, experiment code, and dataset: https://github.com/i-gao/model-equality-testing.
@@ 2 The Model Equality Testing problem
Suppose an auditor is interested in a task parameterized by a distribution $\pi$ over $m$ prompts and a maximum completion length of $L$ tokens. The auditor has sample access to a reference distribution $P$ and API distribution $Q$, both operating on the same vocabulary $\mathcal{V}$ with the same decoding parameters. The auditor samples $N$ prompt-completion pairs [^3] In Appendix C.6, we discuss how to extend this setup to unequal sample sizes. $z:=(x,y)$ from each distribution:
@@ 3 Method
$\widehat{\text{MMD}}(\mathcal{D}_{P},\mathcal{D}_{Q})$ $=\frac{1}{N(N-1)}\left[\sum_{z,z^{\prime}\in\mathcal{D}_{P}}k(z,z^{\prime})+\sum_{z,z^{\prime}\in\mathcal{D}_{Q}}k(z,z^{\prime})\right]-\frac{2}{N^{2}}\sum_{z\in\mathcal{D}_{P}}\sum_{z^{\prime}\in\mathcal{D}_{Q}}k(z,z^{\prime}).$ (4)
@@
The choice of kernel $(k,\phi)$ determines the test’s semantics and power. For example, setting $\phi(z)$ to be an indicator of whether $y$ passes an automated verifier for $x$, when one is available, leads to rejecting the null when $P,Q$ result in different task accuracies. However, because $\phi$ is relatively coarse, the MMD may be zero even when $P\neq Q$, limiting the test’s power. At the other extreme are universal kernels, which guarantee that the MMD is zero if and only if $P=Q$ (Gretton et al., 2012). One such universal kernel for strings is the computationally expensive all-substrings kernel (Borgwardt et al., 2006):
@@
where $y$ shorter than $L$ is right-padded with a special token. Intuitively, a test based on this kernel rejects if a significantly larger number of substitutions are needed to align completions between $\mathcal{D}_{P}$ and $\mathcal{D}_{Q}$ than within each sample. The associated Hamming MMD is a pseudo-metric, as it is zero when $P=Q$ and obeys the triangle inequality, but may not separate all distributions (Appendix A.1). Despite this limitation, we find in the following sections that this kernel is empirically effective and well-suited to common distortions we encounter with language models: quantization, watermarking, finetuning, and related distortions tend to result in detectable inter-sample Hamming distances.
@@ 4.1 Detecting quantization and watermarking
In our first experiments, the reference distribution $P$ represents full-precision weights published on Hugging Face. We evaluate if tests can distinguish $P$ from alternative distributions $Q$:
@@
We repeat evaluations for 5 instruction-tuned models: Mistral 7B Instruct (Jiang et al., 2023), Llama-3 8B and 70B Instruct, and Llama-3.1 8B and 70B Instruct (Meta, 2024). Additional models are in Appendix C.1.
Tests. We compare three choices of kernels: the Hamming kernel (7), the all-substrings kernel (5), and the one-hot kernel (6). We also evaluate two tests from the multinomial testing literature (Balakrishnan & Wasserman, 2018; Bhattacharya & Valiant, 2015):
@@
Figure 2: (Left) Sample complexity of tests. At an average of just 10 samples per prompt, the Hamming MMD test is able to detect quantization and watermarking with nontrivial power. Curves first median power across alternative distributions $Q$, averaged over language models and prompt distributions, with shaded standard errors. Results stratified by language model and alternative are in Appendix C.1. (Middle) While other tests rapidly degrade in power when the user is interested in longer completions, the Hamming MMD test maintains power best across completion lengths. (Right) Power of the Hamming MMD test, stratified by alternative distribution. The test is significantly less powerful against the fp16 alternative.
@@ 4.3 Distinguishing if samples come from different language models
[…]
@@
Llama-3 8B 0.76 (0.00)
Llama-3 70B 1.00 (0.00)
Llama-3.1 70B 0.75 (0.07)
OLMo 7B 1.00 (0.00)
gpt-4o-mini 1.00 (0.00)
Llama-3.1 8B (watermark) 0.32 (0.04)
Llama-3.1 8B (int8) 0.07 (0.01)
Llama-3.1 8B (nf4) 1.00 (0.00)
Figure 3: (Upper left) The Hamming MMD test is able to detect when Llama-3 8B has been finetuned on datasets of $1000$ samples, even after a single epoch. Power is higher, earlier, when the finetuning distribution is i.i.d. with the testing distribution. (Lower left) The Hamming MMD test can also detect when two models are different with near-perfect power. Standard errors are over prompt distributions. Full results are in Appendix C.2. (Right) The MMD framework allows us to estimate statistical distance between any models from which we can draw samples. The cells show average estimated MMDs over $10$ bootstraps. Rows are sorted using spectral clustering with two components. Models within a family are typically clustered together, suggesting that factors like training data, rather than scale, determine model similarity.
Switching to character space makes the testing problem significantly more challenging, as the sample space is higher dimensional: Figure 3 (lower left) shows that power to detect quantization and watermarking is lower in this setting than in the token space setting (§4.1). However, even in this more challenging setting, we find that all model swaps are detectable with 100% power, except for pairs within the Llama family, e.g., Llama-3.1 8B and Llama-3.1 70B (75% power) or Llama-3 8B and Llama-3.1 8B (76% power; Figure 3 lower left). Model swaps are significantly more detectable than watermarking or quantization. Appendix D.2 shows qualitative examples of completions from different models.
@@
A useful feature of the MMD tests is that the test statistic is an estimator of a distance. As a result, we can reuse the machinery to quantify the degree to which two models differ by estimating $\mathbb{E}_{x\sim\pi}[\text{MMD}(P(\cdot\mid x),Q(\cdot\mid x))]$. [^5] Note that the Hamming MMD is a pseudometric: $\text{MMD}_{\text{Hamming}}(P,Q)=0$ does not imply $P=Q$ (Appendix A.1). Figure 3 (right) estimates the Hamming MMD between all pairs of models. To decrease estimator error, we increase the sample size to $N=100m$ samples from each model and report the average $\widehat{\text{MMD}}$ over $10$ simulations, along with standard errors. We observe that models within a family are typically clustered together, suggesting that training data, rather than scale, determines model similarity. Surprisingly, while generations of Llama models (3 and 3.1) are close in distance, some generations of GPT models (e.g., 4-preview and 3.5-turbo) are not. This result suggests the use of different training data or procedures between these models.
@@ 5 Auditing inference API providers
As a case study, we now apply our test to 31 commercial inference endpoints for four of Meta’s Llama models (Figure 1). These endpoints are distributed across nine API providers from Summer 2024: Amazon Bedrock, Anyscale, [^6] We collected samples from Anyscale’s serverless endpoints from before they were deprecated in August 2024. Azure AI Studio, Deepinfra, Fireworks AI, Groq, Perplexity, Replicate, and Together.ai.777 At the time of writing, only two providers disclosed distribution-altering optimizations: Fireworks AI noted semantic caching, and Together.ai noted quantization. We are interested in whether endpoints differ from weights published by Meta on Hugging Face inferenced at commonly accepted precisions. Specifically, we consider two possible null distributions: the full-precision weights ($P_{1}$) and the fp16-precision weights ($P_{2}$). The null and alternative hypotheses are